the sum of three consecutive odd number is 87
find the first number

Answer by mananth(16949)

Let the first number be x
the second and third consequitive odd number will be x+2, and x+4
x+x+2+x+4=87
3x+6=87
3x=87-6
3x=81
x=81/3
x=27
The first number is 27